# From many-body ab initio to effective excitonic models: a versatile mapping approach including environmental embedding effects

### v1

@@ 2.4 Approximate FE-FE and FE-CT exciton couplings
The presence of an embedding environment (e.g. PCM) affects the electrostatic coupling in two ways, as discussed by Mennucci and coworkers. 93, 94 The first trivial effect, results from the change in the transition density of each molecular fragment due to the embedding polarizable medium. The second contribution results from the dielectric screening of the field exerted by the transition densities. Both effects are considered in our calculations, namely, transition densities are obtained from embedded BSE calculations and screened couplings are computed as
where the superscript ”e” labels quantities calculated in the environment and is the medium reaction field defined in Equation 8 that here is calculated at the PCM level. Following Mennucci and coworkers,94, 93 we introduce an effective screening factor
where the screened and unscreened interactions are both computed for the same transition densities.
@@
where is the relative permittivity in the optical range. This expression is derived in the Appendix. To the best of our knowledge, this result has not been reported before. We have explicitly verified the agreement between Equation 20 and numerical calculations based on our PCM implementation.44 For dipoles located at the cavity center, the limiting values are quickly recovered for dipole-dipole distances larger than twice the cavity radius, i.e. once the two dipoles are embedded in separate cavities. We emphasize that two elements concur to the result in Equation 20: (i) the screening of the field of one of the dipoles (source dipole) by the charges induced on the surface of its cavity; (ii) the Onsager cavity field generated within the cavity hosting the probe dipole by the (screened) field of the source dipole.95
The FE-CT and some of the CT-CT couplings can be approximated as the matrix elements of the Kohn-Sham Hamiltonian between frontier molecular orbitals of the involved fragments. Referring to the example in Figure 1 for two-level molecules (HOMO, LUMO), the one-electron approximation for the coupling associated with XHT and HT process between fragments and is , where is the Kohn-Sham Hamiltonian for the dimer - and is HOMO of fragment . Similarly, the matrix element for XET and ET processes reads .
@@ 2.5 Computational details
In this work, we performed PBE0 calculations 96, 97 using the ORCA package 98 to obtain the Kohn-Sham wavefunctions. and BSE calculations were carried out with the BeDEFT program.99 All calculations were done with the def2-TZVP basis set 100 a the corresponding def2-TZVP-RI auxiliary basis for Coulomb fitting in the resolution-of-the-identity framework (RI-V)101. ev calculations have been performed, correcting 4, 8, and 12 states in calculations of single fragments, dimers, and trimers, respectively. For PCM embedding, we adopted a dielectric constant , typical of non-polar organic solids. Other parameters were set to the default values for benzene from the Minnesota Solvent Descriptor Database. DFT calculations employed the ORCA-code implementation of the conductor-like polarizable continuum model (C-PCM).102 A double-layer formulation of the integral equation formalism PCM model (IEF-PCM), allowing for charge-spilling effects, has been used in and BSE computations.44 The ground-state embedding of calculations on the pentacene crystal employed the charge response model88 as implemented in the MESCal code.89 The pentacene molecular geometry employed in model dimer calculations (Section 3.1) has been optimized at the B3LYP/def2-TZVP level in the gas phase. Dimers have been then built by creating a translated replica of the molecule along the direction perpendicular to the molecular plane, obtaining parallel co-facial dimers for different intermolecular distances. Calculations on the pentacene crystal (Section 3.2) have been performed for the Siegriest triclinic structure103.
@@ 3.1.1 Pentacene homodimer in vacuum
which upon diagonalization yields eigenvalues that reproduce the 4 lowest-energy singlet BSE excitation energies (i.e. 1.870, 1.920, 2.924, and 2.930 eV) within 1 meV tolerance. The diagonal elements in Eq. LABEL:e:Heff2 represent the diabatic energies that are plotted for different inter-fragment distances in Fig. 2a, together with the BSE energies and the ev HOMO-LUMO gap. We observe that the diabatic energies follow the behavior expected from fundamental physical considerations. The diabatic energies of the two localized FEs remain constant with the inter-fragment distances, rapidly approaching the two degenerate FEs of the dimer in the limit. The diabatic energy of CT states follows the behavior prescribed by the e-h interaction for these intermolecular excitons, converging toward the photoemission gap at large . The diabatic energies converge to the adiabatic dimer excitation energies for , while excitonic couplings determine significant deviations at distances typical of intermolecular contacts.
@@ 3.1.2 Pentacene homodimer in PCM
The FE-FE coupling (Figure 3b) obtained with the diabatization scheme closely matches the value computed as screened interaction between transition densities (Equation 18), recovering the screened dipolar approximation at large . The coupling between FEs in PCM is significantly reduced by screening phenomena as compared to gas-phase values. Figure 4 shows the effective screening factor, corresponding to the ratio between screened and unscreened Coulomb interaction between transition densities (see Equation 19), as a function of . The screening factor decays monotonically from 0.7 at 4 Å distance to reach an asymptotic value of 0.55 at large . The trend is similar to what was reported by Mennucci and coworkers for a large set of chromophores in PCM.93, 94 We further note that the limiting values is within 12% from the analytical result for point dipoles in spherical cavities (Equation 20), i.e. for (horizontal line in Figure 4). Such reasonable agreement shall be considered accidental, since for anisotropic cavities, the screening factor largely depends on the dipole orientation (see SI, Figure S1).
@@ 3.2 Pentacene molecular crystal
We now apply our projective diabatization scheme to the pentacene crystal, which allows us to discuss some important and often overlooked aspects of the ab initio calculation of the parameters entering model Hamiltonians for extended systems. Before proceeding, we recall two technical aspects of our calculations that will be crucial for the quality of the results presented in the following. First, all calculations in the crystal employed a QM/MM scheme employing a polarizable atomistic model in the ground-state DFT calculation, where the chosen QM region (single molecule, dimer, etc.) is embedded in the electrostatic potential of the bulk crystal. The environmental screening of electronic excitations at the and BSE level has been described at the PCM level (, see Section 2.2 for details). The proper account of the electrostatic landscape and dielectric screening of the environment is essential for the reliable transfer of the parameters calculated from a small QM subsystem to crystals. In the second instance, the definition of the localized (diabatic) basis functions relies on a single (DFT, , BSE) calculation for each of the symmetry-unique molecules in the unit cell. Translational invariance is enforced by creating translational copies of unique fragments and defining the basis functions accordingly. This ensures phase consistency among all basis FE and CT states throughout a virtually infinite crystal, hence determining unambiguously the signs of the off-diagonal elements in model Hamiltonians.104, 3
@@
Trimer calculations have been thus performed to include these effects, specifically computing the and systems as defined in Figure 5(a)c,d. The construction of the diabatic basis for trimers follows the same prescription adopted for dimers, including 3 FEs and 6 CT states, resulting in a 9-state model. The values of all matrix elements are reported in the SI, the following discussion focuses on the most relevant ones.
As expected, ET and HT couplings are sizeable, with a magnitude of tens of meV. The comparison of the Hamiltonian obtained for the two trimers reveals that the et coupling is almost independent of the position of the hole, namely meV and meV. Similarly, also HT is almost unaffected by the electron position, i.e. meV and meV. The fact that these hopping terms barely depend on the position of the geminate charge, indicates that the screened-Coulomb () and exchange terms in the BSE Hamiltonian (see Equation 4) are almost negligible, as somewhat expected for loosely overlapping orbitals localized on different molecules. We note that these two HT and ET couplings are similar to their one-body approximation, meV and meV for HT and ET, respectively. However, the quantitative discrepancy between these Kohn-Sham and BSE/ couplings attests to the importance of many-body effects for these interactions.
@@
The diagonal energies of CT states are the parameters that are mostly sensitive to the environment, as they consist of localized charges (in contrast with FE that are neutral species) that directly probe the local electrostatic potential within the crystal, and that experience the response of the polarizable medium to their strong dipole field. The largest variability is found for the energy of CT1→3, i.e. 47 meV, below 2% of the excitation energy. This is a very positive result, being this variability smaller than the accuracy of the BSE/-based QM/MM methodology (typically 0.1-0.2 eV). For comparison, calculation not accounting for any embedding (i.e. computing dimers and trimers in the gas phase), yields CT states whose energy varies in the range of 0.48 eV, one order of magnitude larger than embedded calculations. We note that most of the spread is cured by the electrostatic embedding in the ground state, which compensates for the otherwise different electrostatic fields at molecular sites developing in gas-phase dimers and trimers. Polarization effects on excitations further reduce the spread in CT states energies obtained from calculations on different QM subsystems and, most importantly, are crucial to reproduce excitation energies in a condensed phase correctly. The spread in the values obtained for off-diagonal matrix elements is found to be smaller than 1 meV, as shown in Table 2.
@@ 4 Conclusions
G.D. acknowledges the contribution from Jing Li in the first steps of this research and thanks Samuele Giannini for stimulating discussions. This work received financial support from the French ”Agence Nationale de la Recherche”, project RAPTORS (ANR-21-CE24-0004-01). This work was performed using HPC resources from GENCI-TGCC (Grant No. 2021-A0110910016).
Appendix A Appendix: Interaction of two dipoles inside separated spherical cavities
A.1 The cavity response to an interior point multipole
In the following, we consider a continuous polarizable medium with a dielectric constant . Let us start placing the center of a cavity of radius as the origin and setting the position of a point charge at . This corresponds to a standard electrostatic problem which is solved by considering first that in the absence of the cavity, the potential due to the point charge can be expressed using Legendre polynomials of integer order
with being the vacuum permittivity. In our case, we simply generalize this definition to (non-normed) point multipoles by differentiating the formulas with respect to
In particular, the term defines the potential and normal field due to the multipole in the long-range limit
Similarly, the reaction field from the cavity can be written in the general form
with the derivatives along
The coefficients can then be evaluated by noting that the radial component of the total electric displacement field is continuous across the cavity boundary at r = a
which leads in particular to (sorting out the factors)
Thence, the resulting general long-range limit for the renormalization of the potential outside a spherical cavity, for a multipole placed within the cavity:
Introducing the relative permittivity , we recover the usual charge potential reaction field renormalization for ,
while for we recover the corresponding expression for a dipole distribution
A.2 The cavity response to an exterior field
We consider now a homogeneous electric field (along the x coordinate) so that the potential reads
(here we simply used the fact that ). Once again, we write the reaction field from the cavity and its derivative using equations (29) and (30) and evaluate the coefficients by noting that the radial component of the total electric displacement field is continuous across the boundary r = a
which brings immediately
with the corresponding cavity reaction potential
so that the total potential seen from within the cavity is
Here, we get the relation between the total electric field seen from within the cavity and the applied external field
which is the Onsager result for the cavity field.95
A.3 The dipole-dipole interaction across separate cavities
In this case, we combine results from the previous sections. The dipoles within cavity 1 centered in generates a screened potential outside its cavity that was derived in the previous sections
The resulting electric field reads
so that in the vicinity of the second cavity placed at the origin , the potential can be expanded as
with
and
Applying the result (40) from the previous section, we get that the total potential , as seen from a charge distribution placed within the second cavity reads
A charge placed in the second cavity will thus probe the first term of (47) right-hand side, leading to a charge-dipole interaction renormalized by
while a dipole placed in the second cavity will only probe the second term of (47) right-hand side, leading to a dipole-dipole interaction renormalized by

### v2

@@ 2.4 Approximate excitonic couplings
2.5 Dielectric screening of FE-FE couplings
The presence of a dielectric environment (e.g. PCM) largely affects FE-FE couplings. Restricting ourselves to the leading electrostatic interaction between transition densities (Eq. 16), the embedding medium affects the coupling in two ways, as discussed by Mennucci and coworkers. 100, 101 The first trivial effect, results from the change in the transition density of each molecular fragment due to the embedding polarizable medium. The second contribution is due to the dielectric screening of the field exerted by the transition densities. Both effects are considered in our calculations, namely, transition densities are obtained from embedded BSE calculations and screened couplings are computed as
where the superscript “e” labels quantities calculated in the environment and is the medium reaction field defined in Equation 8 that here is calculated at the PCM level. Following Mennucci and coworkers,101, 100 we introduce an effective screening factor
where the screened and unscreened interactions are both calculated for the same transition densities, here obtained in PCM. Computing the ratio for the same transition density is important to properly disentangle the two effects of the polarizable medium on the excitonic coupling, namely the impact of the environment on the transition density and the dielectric screening of the interaction.
@@ 2.5 Dielectric screening of FE-FE couplings
where is the relative permittivity in the optical range. The derivation of this expression is provided as Supporting Information. To the best of our knowledge, this result has not been reported before. We have explicitly verified the agreement between Equation 20 and numerical calculations based on our PCM implementation.50 For dipoles located at the cavity center, the limiting values are quickly recovered for dipole-dipole distances larger than twice the cavity radius, i.e. once the two dipoles are embedded in separate cavities. We emphasize that two elements concur to the result in Equation 20: (i) the screening of the field of one of the dipoles (source dipole) by the charges induced on the surface of its cavity; (ii) the Onsager cavity field generated within the cavity hosting the probe dipole by the (screened) field of the source dipole.102
@@ 2.6 Computational details
In this work, we performed PBE0 calculations 103, 104 using the ORCA package 105 to obtain the Kohn-Sham wavefunctions. and BSE calculations were carried out with the BeDEFT program.106 All calculations were done with the def2-TZVP basis set 107 a the corresponding def2-TZVP-RI auxiliary basis for Coulomb fitting108 in the resolution-of-the-identity framework (RI-V).35 ev calculations have been performed, correcting 4, 8, and 12 states in calculations of single fragments, dimers, and trimers, respectively. For PCM embedding, we adopted a dielectric constant , typical of non-polar organic solids. Other parameters were set to the default values for benzene from the Minnesota Solvent Descriptor Database. DFT calculations employed the ORCA-code implementation of the conductor-like polarizable continuum model (C-PCM).109 A double-layer formulation of the integral equation formalism PCM model (IEF-PCM), allowing for charge-spilling effects, has been used in and BSE computations.50 The ground-state embedding of calculations on the pentacene crystal employed the charge response model95 as implemented in the MESCal code.96 The pentacene molecular geometry employed in model dimer calculations (Section 3.1) has been optimized at the B3LYP/def2-TZVP level in the gas phase. Dimers have been then built by creating a translated replica of the molecule along the direction perpendicular to the molecular plane, obtaining parallel co-facial dimers for different intermolecular distances. Calculations on the pentacene crystal (Section 3.2) have been performed for the Siegriest triclinic structure110.
@@ 3.1.1 Pentacene homodimer in vacuum
which upon diagonalization yields eigenvalues that reproduce the 4 lowest-energy singlet BSE excitation energies (i.e. 1.870, 1.920, 2.924, and 2.930 eV) within 1 meV tolerance. The diagonal elements in Eq. 21 represent the diabatic energies that are plotted for different inter-fragment distances in Fig. 2a, together with the BSE energies and the ev HOMO-LUMO gap. We observe that the diabatic energies follow the behavior expected from fundamental physical considerations. The diabatic energies of the two localized FEs remain constant with the inter-fragment distances, rapidly approaching the two degenerate FEs of the dimer in the limit. The diabatic energy of CT states follows the behavior prescribed by the e-h interaction for these intermolecular excitons, converging toward the photoemission gap at large . The diabatic energies converge to the adiabatic dimer excitation energies for , while excitonic couplings determine significant deviations at distances typical of intermolecular contacts.
@@ 3.1.2 Pentacene homodimer in PCM
The FE-FE coupling (Figure 3b) obtained with the diabatization scheme closely matches the value computed as screened electrostatic interaction between transition densities (Equation 18), recovering the screened dipolar approximation at large . The Coulomb coupling between the transition densities of the two FEs in PCM is significantly reduced as compared to the unscreened interaction, as shown in Figure 4a. We recall the importance of computing screened and unscreened interactions between the same transition densities, here corresponding to those obtained from BSE/ calculations in PCM. This is crucial for disentangling the effect of dielectric screening from the effect of the polarizable environment on the transition density. Figure 4b shows the effective screening factor, corresponding to the ratio between screened and unscreened Coulomb interaction between transition densities (see Eq. 19), as a function of . The screening factor decays monotonically from 0.76 at 3.5 Å distance to reach an asymptotic value of 0.55 at large . The trend is similar to what was reported by Mennucci and coworkers for a large set of chromophores in PCM.100, 101 We further note that the limiting values is within 12% from the analytical result for point dipoles in spherical cavities (Eq. 20), i.e. for (horizontal line in Fig. 4a). Such reasonable agreement shall be considered accidental, since for anisotropic cavities, the screening factor largely depends on the dipole orientation (see SI, Figure S1).
@@ 3.2 Pentacene molecular crystal
We now apply our projective diabatization scheme to the pentacene crystal, which allows us to discuss some important and often overlooked aspects of the ab initio calculation of the parameters entering model Hamiltonians for extended systems. Before proceeding, we recall two technical aspects of our calculations that will be crucial for the quality of the results presented in the following. First, all calculations in the crystal employed a QM/MM scheme based on a polarizable atomistic model in the ground-state DFT calculation, where the chosen QM region (single molecule, dimer, etc.) is embedded in the electrostatic potential of the bulk crystal. The environmental screening of electronic excitations at the and BSE level has been described at the PCM level (, see Section 2.2 for details). The proper account of the electrostatic landscape and dielectric screening of the environment is essential for the reliable transfer of the parameters calculated from a small QM subsystem to crystals. In the second instance, the definition of the localized (diabatic) basis functions relies on a single (DFT, , BSE) calculation for each of the symmetry-unique molecules in the unit cell. Translational invariance is enforced by creating translational copies of unique fragments and defining the basis functions accordingly. This ensures phase consistency among all basis FE and CT states throughout a virtually infinite crystal, hence determining unambiguously the signs of the off-diagonal elements in model Hamiltonians.111, 3
@@
Trimer calculations have been thus performed to include these effects, specifically computing the and systems as defined in Figure 5(a)c,d. The construction of the diabatic basis for trimers follows the same prescription adopted for dimers, including 3 FEs and 6 CT states, resulting in a nine-state model. The upper triangular part of the Hamiltonians (in eV) for the two trimers reads:
The ET and HT couplings, missing in dimer calculations, are sizeable, with a magnitude of tens of meV. The comparison of the Hamiltonian obtained for the two trimers reveals that the ET coupling is almost independent of the position of the hole, namely meV and meV. Similarly, also HT is almost unaffected by the electron position, i.e. meV and meV. The fact that these hopping terms barely depend on the position of the geminate charge, indicates that the screened-Coulomb () and exchange terms in the BSE Hamiltonian (see Equation 4) are almost negligible, as somewhat expected for loosely overlapping orbitals localized on different molecules. We note that these two HT and ET couplings are similar to their one-body approximation, meV and meV for HT and ET, respectively. However, the quantitative discrepancy between these Kohn-Sham and BSE/ couplings attests to the importance of many-body effects for these interactions.
@@
The diagonal energies of CT states are the parameters that are mostly sensitive to the environment, as they consist of localized charges (in contrast with FE that are neutral species) that directly probe the local electrostatic potential within the crystal, and that experience the response of the polarizable medium to their strong dipole field. The largest variability is found for the energy of CT1→3, i.e. 47 meV, below 2% of the excitation energy. This is a very positive result, being this variability smaller than the accuracy of the BSE/-based QM/MM methodology (typically 0.1-0.2 eV). For comparison, calculation not accounting for any embedding (i.e. computing dimers and trimers in the gas phase), yields CT states whose energy varies in the range of 0.48 eV, one order of magnitude larger than embedded calculations. We note that most of the spread is cured by the electrostatic embedding in the ground state, which compensates for the otherwise different electrostatic fields at molecular sites developing in gas-phase dimers and trimers. Polarization effects on excitations further reduce the spread in CT states energies obtained from calculations on different QM subsystems and, most importantly, are crucial to quantitatively reproduce excitation energies in a condensed phase. The spread in the values obtained for off-diagonal matrix elements is found to be smaller than 1 meV, as shown in Table 1.
@@ 4 Conclusions
G.D. acknowledges the contribution of Jing Li in the first steps of this research and thanks Samuele Giannini for stimulating discussions. This work received financial support from the French “Agence Nationale de la Recherche”, project RAPTORS (ANR-21-CE24-0004-01). This work was performed using HPC resources from GENCI-TGCC (Grant No. 2021-A0110910016).
Explicit expression for a Frenkel-CT excitonic model Hamiltonian. Supplementary calculation results. Derivation of Eq. 20 for the screened interaction between point dipoles in spherical cavities and additional results for screened dipolar interaction in realistic molecular cavities. Bartels-Stewart algorithm for the solution of the Sylvester’s equation.